OpenStudy (atlchic):

Solve 3/4x-4=7+1/2x. Show your work.

OpenStudy (anonymous):

x = 1/44

OpenStudy (atlchic):

show work plz

OpenStudy (atlchic):

show work plz

OpenStudy (anonymous):

-(16 x-3)/(4 x) = (14 x+1)/(2 x)(2/2) -(16 x-3)=(14 x+1) x=1/44

OpenStudy (atlchic):

thks, i have another question

OpenStudy (precal):

close this one and post a new problem (please)